On-time arrivals, lost baggage, and customer complaints are three measures that are typically used to measure the quality of service being offered by airlines. Suppose that the following values represent the on-time arrival percentage, amount of lost baggage, and customer complaints for 10 U.S. airlines.



Airline

On-Time
Arrivals (%)

Mishandled Baggage
per 1,000 Passengers
Customer
Complaints per
1,000 Passengers
Virgin America 83.5 0.87 1.50
JetBlue 79.1 1.88 0.79
AirTran Airways 87.1 1.58 0.91
Delta Air Lines 86.5 2.10 0.73
Alaska Airlines 87.5 2.93 0.51
Frontier Airlines 77.9 2.22 1.05
Southwest Airlines 83.1 3.08 0.25
US Airways 85.9 2.14 1.74
American Airlines 76.9 2.92 1.80
United Airlines 77.4 3.87 4.24
(a) Based on the data above, if you randomly choose a Delta Air Lines flight, what is the probability that this individual flight will have an on-time arrival? If required, round your answer to three decimal places.
(b) If you randomly choose 1 of the 10 airlines for a follow-up study on airline quality ratings, what is the probability that you will choose an airline with less than two mishandled baggage reports per 1,000 passengers? If required, round your answer to two decimal places.
(c) If you randomly choose 1 of the 10 airlines for a follow-up study on airline quality ratings, what is the probability that you will choose an airline with more than one customer complaint per 1,000 passengers? If required, round your answer to two decimal places.
(d) What is the probability that a randomly selected AirTran Airways flight will not arrive on time? If required, round your answer to three decimal places.
Answer #1

a)

probability that this individual flight will have an on-time arrival =0.865 (since it has on time arrival of 86.5%)

b)

probability =3/10 =0.3 (since 3 of 10 airlines have less than two mishandled baggage reports per 1,000 )

c)

probability =5/10 =0.5 (since 5 of 10 airlines have more than one customer complaint per 1,000 passengers

d)

probability that a randomly selected AirTran Airways flight will not arrive on time =1-P(arrive on time) =1-0.871 =0.129